Sea water has sodium ion concentration of 1.08 x 104 ppm. If the sodium is present in the form of dissolved NaCl, what mass of NaCl is in each liter of sea water? density of sea water is 1.05 g/mL
1
Expert's answer
2017-11-14T09:16:07-0500
1.08×104 ppm Na = 1.08×104 mg Na/1000 g solution
Density 1.05 g/ml = 1050 g/L
1.08×104 mg Na/1000 g solution ×1005 g/L= 1.13×104 mg Na/L
1.13×104 mg Na/L ×1 g/1000 mg = 11.3 g Na/L
M(Na)=23 g/mol
M(NaCl)=58.44 g/mol
m(NaCl)=11.3 g Na/L×58.44 g NaCl/23.0 g Na=28.7 g
Answer
Mass of NaCl must be 28.7 g in each liter of sea water.